# Rigidity of anti-de Sitter (2+1)-spacetimes with convex boundary near the Fuchsian locus

### v1

@@ 3.1. Equivariant surfaces in Minkowski 3-space
[…]
Under the identification above between operator fields and Riemannian metrics, an operator field $a$ is orthogonal to the tangent space to the isotopy orbit of $\sigma$ if and only if it is Codazzi for $\sigma$. See, e.g., [65, Section 3.2]. It is also easy to notice that it is orthogonal to the tangent space to the conformal orbit of $\sigma$ if and only if it is traceless. We arrive to the well-known fact that every tangent vector to the Teichmüller space can be represented by a $\sigma$-symmetric traceless Codazzi operator (which is then unique). When $\sigma$ is the hyperbolic representative of the conformal class, the $L^{2}$-scalar product produces the Weil–Petersson scalar product on the Teichmüller space, up to a universal constant depending on a convention. It also follows from the discussion that if $x_{1},x_{2}\in T_{\sigma}\mathscr{T}$ and they are represented by symmetric Codazzi fields $a_{1},a_{2}$ such that $a_{2}$ is traceless, then
[…]
Now return to the equivariant isometric immersion $(\rho,f)$ of $h$. Define
[…]
Due to Theorem 3.2, this is indeed a well-defined function of a negatively curved metric $s$ and a hyperbolic metric $\sigma$. Furthermore, it depends only on the isotopy classes of $s$, $\sigma$. For a fixed $s$, we consider then $F(s,.)$ as a function $F_{s}$ over the Teichmüller space $\mathscr{T}$. It was shown by Smith in [58, Lemma 5.3] that
@@
If a tangent vector in $T_{\sigma}\mathscr{T}$ is represented by a traceless $\sigma$-symmetric Codazzi field $a^{\prime}$, then
[…]
@@
The vector field $-x^{\prime}_{s}$ is the Weil–Petersson gradient field of $F_{s}$ and $-x_{s}$ is the symplectic Weil–Petersson gradient field of $F_{s}$.
Furthermore, Smith proved [58, Lemma 5.6]
@@
For any two negatively curved metrics $s_{+}$ and $s_{-}$ the vector fields $x_{s_{+}}^{\prime}$ and $-x_{s_{-}}^{\prime}$ intersect transversely at every intersection point as submanifolds of $T\mathscr{T}$. The same holds for $x_{s_{+}}$ and $-x_{s_{-}}$.
@@
If at a point $\sigma\in\mathscr{T}$ we have $x_{s_{+}}^{\prime}(\sigma)=-x_{s_{-}}^{\prime}(\sigma)$, then the Weil–Petersson gradients of functions $F_{s_{+}}$ and $-F_{s_{-}}$ coincide at $\sigma$. Thus, $\sigma$ is a critical point of $F_{s_{+}}+F_{s_{-}}$. Since the latter function is strictly convex with respect to the Weil–Petersson metric, its Weil–Petersson Hessian is non-degenerate. Thereby, the intersection point of $x_{s_{+}}^{\prime}$ and $-x_{s_{-}}^{\prime}$ is transverse at $x_{s_{+}}^{\prime}(\sigma)=-x_{s_{-}}^{\prime}(\sigma)$. The same holds for the vector fields $x_{s_{+}}$, $-x_{s_{-}}$. ∎
@@
Let $\rho:\pi_{1}S\rightarrow G^{0}$ be a quasi-Fuchsian representation and $f_{+},f_{-}:\widetilde{S}\rightarrow\mathbb{R}^{2,1}$ be two $\rho$-equivariant strictly convex embeddings, one is future-convex and one is past-convex. Let $\dot{\rho}\in Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g}^{0})$ be a cocycle and $v$ be a $\dot{\rho}$-automorphic isometric vector field on $f_{+}(\widetilde{S})$ and $f_{-}(\widetilde{S})$. Then $v$ is trivial.
@@
Let $s_{+}$ and $s_{-}$ be the induced metrics, $b_{+}$ and $b_{-}$ be the shape operators of $f_{+}$ and $f_{-}$ and $\dot{s}_{+}$ and $\dot{s}_{-}$ be their variations induced by $v$. If $\dot{b}_{+}=0$, then $v$ is trivial on $f_{+}(\widetilde{S})$. Hence $\dot{\rho}$ is a coboundary. Then, due to Theorem 3.2, also $\dot{b}_{-}=0$ and hence $v$ is trivial also on $f_{-}(\widetilde{S})$. Furthermore, on both surfaces it is the restriction of the same Killing field, determined by $\dot{\rho}$. Thus, we can assume that both $\dot{b}_{+}$ and $\dot{b}_{-}$ are nonzero. By definition, we have $\phi^{0}_{s_{+},+}(b_{+})=\phi^{0}_{s_{-},-}(b_{-})$. Since the restrictions of $v$ to both $f_{+}(\widetilde{S})$ and $f_{-}(\widetilde{S})$ are $\dot{\rho}$-automorphic, by construction,
@@
Due to Theorem 3.2, expression (3.3) is nonzero. But this contradicts Corollary 3.6, as the images of ϕs+,+subscriptitalic-$\phi_{s_{+},+}$ and ϕs,subscriptitalic-$\phi_{s_{-},-}$ are exactly the vector fields $x_{s_{+}}$ and $-x_{s_{-}}$, when the latter are considered as submanifolds of $T\mathscr{T}$. ∎
@@ 3.2. The infinitesimal Pogorelov map
Recall the definition of $\mathbb{A}d\mathbb{S}^{3}$ from Section 2.1. We now pick a convenient affine chart for it, given by a projection to the plane $\{x\in\mathbb{R}^{2,2}:x_{4}=1\}$. This plane is also a model for Minkowski 3-space, and we will consider it endowed simultaneously with both geometries. We denote it now by $\mathbb{R}^{2,1}$, and by the origin we will mean the point $o=(0,0,0,1)$. The part of $\mathbb{A}d\mathbb{S}^{3}$ that is captured in this plane is the interior of a one-sheeted hyperboloid
@@
Here, for convenience, we only consider the restrictions of the Killing fields to $C$, though one can define $\Phi$ over the whole $TA$, see, e.g., [20]. See [24, Lemma 11] of Fillastre for a proof of Lemma 3.8, which is based on [54, Lemma 1.10] of Schlenker. In particular, $\Phi$ determines an isomorphism $\Psi:\mathfrak{g}\rightarrow\mathfrak{g}^{0}$ of Lie algebras, considered as vector spaces. We can specify further
@@
Proof.
We have decompositions $\mathfrak{g}=\mathfrak{g}_{r}\oplus\mathfrak{g}_{t}$, $\mathfrak{g}^{0}=\mathfrak{g}_{r}^{0}\oplus\mathfrak{g}_{t}^{0}$ into pure rotations and pure translations with respect to $o$. Here also $\mathfrak{g}_{r}\cong\mathfrak{g}_{r}^{0}\cong\mathfrak{g}_{F}$ via the chosen inclusions of $G_{F}$ to $G$ and $G^{0}$. Killing fields from the first summands are characterized by the property that they do not have radial component. Killing fields from the second summands are characterized by the property that along some geodesic passing through $o$ they do not have a lateral component. (Here the geodesic depends on a field.) It is clear from this description that both $G_{F}$ and $\Psi$ preserve these decompositions. Furthermore, $\Psi$ preserves the isomorphisms $\mathfrak{g}_{r}\cong\mathfrak{g}_{r}^{0}\cong\mathfrak{g}_{F}$.
The $G_{F}$-action on the first components coincides with the $G_{F}$-action on $\mathfrak{g}_{F}$ and is preserved by $\Psi$. It is an easy observation that for $\gamma\in G_{F}$ and $\xi\in\mathfrak{g}_{t}$ that is an infinitesimal translation along a line $l$ passing through $o$, $\gamma\xi$ is an infinitesimal translation along the line $\gamma l$ with the same norm. It follows that $\Psi$ preserves the $G_{F}$-actions on $\mathfrak{g}$ and $\mathfrak{g}^{0}$. ∎
For a Fuchsian $\rho:\pi_{1}\rightarrow G_{F}$, Lemma 3.9 establishes a canonical isomorphism
$\Psi_{\rho}:Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g})\rightarrow Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g}^{0}),$
which sends coboundaries onto coboundaries, and hence also induces an isomorphism of cohomologies.
Now we also note
@@
Let $\rho\in\widehat{\mathcal{R}}_{F}$, $\Sigma\subset C$ be a $\rho$-invariant submanifold and $v$ be a vector field over $\Sigma$, automorphic with respect to $\dot{\rho}\in Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g})$. Then $\Phi(v)$ is automorphic with respect to $\Psi_{\rho}(\dot{\rho})\in Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g}^{0})$.
@@
$d\rho(\gamma).\big{(}\Phi(v(x))+\Psi_{\rho}(\dot{\rho})(\gamma)(x)\big{)}=d\rho(\gamma).\left(\frac{v_{r}(x)+\dot{\rho}(\gamma)_{r}(x)}{1-r^{2}(x)}+v_{l}(x)+\dot{\rho}(\gamma)_{l}(x)\right)=$
$=\frac{v_{r}(\rho(\gamma)x)}{1-r^{2}(x)}+v_{l}(\rho(\gamma)x)=\Phi(v(\rho(\gamma)x)),$
@@
Proof of Lemma 3.1..
Consider $e\in\mathcal{E}_{F}$. Lift it to a triple $(\rho,f_{+},f_{-})\in\widehat{\mathcal{E}}_{F}$ of a representation $\rho:\pi_{1}S\rightarrow G_{F}$, where $G_{F}$ is considered as a subgroup of $G$, and of $\rho$-equivariant developing maps $f_{+},f_{-}:\widetilde{S}\rightarrow\mathbb{A}d\mathbb{S}^{3}$. We assume that $o$ is the fixed point of $G_{F}$ as a subgroup of $G$. Then the images of $f_{+}$ and $f_{-}$ are contained in $C$. Pick $\dot{e}\in T_{e}\mathcal{E}$ such that $d_{e}\mathcal{I}(\dot{e})=0$. It produces a cocycle $\dot{\rho}\in Z^{1}_{\rho}(\pi_{1}S,\mathfrak{g})$ and a $\dot{\rho}$-automorphic vector field $v$ on $f_{+}(\widetilde{S})$ and $f_{-}(\widetilde{S})$, isometric with respect to the anti-de Sitter metric.
Apply the Pogorelov map $\Phi$ to $v$. The combination of Lemma 3.11 and Lemma 3.10 implies that $\Phi(v)$ is a $\Psi_{\rho}(\dot{\rho})$-automorphic vector field on $f_{+}(\widetilde{S})$ and $f_{-}(\widetilde{S})$, isometric with respect to the Minkowski metric. Hence, by Lemma 3.7, $\Phi(v)$ is trivial, thereby so is $v$. It follows that $\dot{e}=0$. ∎
@@
It is interesting to notice that even though one can define the Pogorelov map outside of $C$, this proof works only when the pair of embeddings is Fuchsian. The reason is that for a general equivariant embedding, its image in the projective model will not be equivariant with respect to some Minkowski representation. There is a nice way to do a rescaling, to obtain quasi-Fuchsian Minkowski spacetimes from anti-de Sitter ones (more precisely, from the components of the complements to the convex core), well-behaved on holonomies, see [7] by Benedetti–Bonsante. However, this is not a projective map, and hence does not have a related Pogorelov map. (A Pogorelov map can be associated to any projective map between pseudo-Riemannian manifolds, see [26]).
@@ 4.1. A bit of functional analysis
For a bounded operator $\mathcal{L}$ between two Banach spaces $\mathcal{A},\mathcal{B}$ we define its reduced minimum modulus
$\gamma(\mathcal{L}):=\inf\{\|\mathcal{L}a\|:a\in\mathcal{A},d(a,\ker\mathcal{L})=1\}.$
When the range of $\mathcal{L}$ is closed, $\gamma(\mathcal{L})>0$. See, e.g., [36, Chapter IV, Theorem 5.2].
For two subspaces $A,B$ of a Banach space we define their distance as
@@
Let $A$ be a finite-dimensional subspace in a Banach space, and $A_{i}$ be a sequence of subspaces, converging to $A$ in the distance (4.1). Let $a_{i}\in A_{i}$ be a sequence of unit vectors. Then, up to passing to a subsequence, $a_{i}$ converges to $a\in A$.
@@
Since the span of the union of $A$ with any vector is finite-dimensional, for every $a_{i}$ there is $a^{\prime}_{i}$ realizing the distance from $a_{i}$ to $A$. Up to passing to a subsequence, all $a^{\prime}_{i}$ belong to the closed ball in $A$ of radius 2. Since it is compact, a subsequence of $a^{\prime}_{i}$ converges to some $a$. It is easy to see that the same subsequence of $a_{i}$ also converges to $a$. ∎
@@ 4.2. Local infinitesimal rigidity
Lemma 4.2.
Let $e\in\mathcal{E}_{F}$. There exists its neighborhood $U_{e}$ in $\mathcal{E}$ such that for every $e^{\prime}\in U_{e}$ the differential $d_{e^{\prime}}\mathcal{I}$ is injective.
@@
Consider $e^{\prime}\in\mathcal{E}$. Let $s_{+}$, $s_{-}$, $b_{+}$ and $b_{-}$ be the respective induced metrics and shape operators on $S$ given by $e^{\prime}$. The differential $d_{e^{\prime}}\mathcal{I}$ is injective if and only if the images of ϕs+,+subscriptitalic-$\phi_{s_{+},+}$ and ϕs,subscriptitalic-$\phi_{s_{-},-}$ intersect transversely at $\phi_{s_{+},+}(b_{+})=\phi_{s_{-},-}(b_{-})$.
@@
Fix a lift $(\rho,f_{+},f_{-})\in\widehat{\mathcal{E}}$ of $e^{\prime}$. Pick $\dot{e}\in T_{e^{\prime}}\mathcal{E}$. If $d_{e^{\prime}}\mathcal{I}(\dot{e})=0$, then $\dot{e}$ can be represented by $(\dot{\rho},\dot{f}_{+},\dot{f}_{-})$ such that $\dot{f}_{+}$ and $\dot{f}_{-}$ are isometric on $f_{+}(\widetilde{S})$ and $f_{-}(\widetilde{S})$. Suppose that $\dot{f}_{+}$ induces zero variation on the shape operator of $f_{+}(\widetilde{S})$. Then $\dot{f}_{+}$ is trivial. Then $\dot{\rho}$ is a coboundary and $\dot{e}=0$. Suppose now that it is not the case and $\dot{b}_{+}$ and $\dot{b}_{-}$ are nonzero variations of the shape operators. Then, by construction, $d_{b_{+}}\phi_{s_{+},+}(\dot{b}_{+})=d_{b_{-}}\phi_{s_{-},-}(\dot{b}_{-})$, which is nonzero due to Theorem 2.2. Thus ϕs+,+subscriptitalic-$\phi_{s_{+},+}$ and ϕs,subscriptitalic-$\phi_{s_{-},-}$ do not intersect transversely. The converse is shown in the same way. ∎
Now we need to show that if $s_{i}$ converge to $s$ and $\rho_{i}\in{\rm im}(\phi_{s_{i},\pm})$ converge to $\rho\in{\rm im}(\phi_{s,\pm})$, then the tangent space to ${\rm im}(\phi_{s_{i},\pm})$ at $\rho_{i}$ converge to the tangent space to ${\rm im}(\phi_{s,\pm})$ at $\rho$. Let $\mathcal{A}$ be the Sobolev space $L^{2}_{1}(T^{1}_{1}S)$ of sections of $T^{1}_{1}S$ that have finite $L^{2}$-norm and so do their first weak derivatives. We consider $\mathcal{A}$ as a Banach space. The desired convergence result follows from
@@
For every $i$ denote by $G_{i}\subset T^{1}_{1}S$ the subbundle of $s_{i}$-symmetric operator fields $TS\rightarrow TS$ satisfying the Gauss equation for $s_{i}$. Denote by $J\subset T^{1}_{1}S$ the subbundle of almost-complex structures on $S$ (compatible with the orientation) and let $j_{i}$, $j$ be the almost-complex structures of $s_{i}$, $s$. Consider the bundle isomorphisms $K_{i}:T^{1}_{1}S\rightarrow T^{1}_{1}S$ given by
@@
where $\lambda_{i}:=\sqrt{-\kappa_{i}-1}$ for the curvature $\kappa_{i}$ of $s_{i}$. Notice that $K_{i}$ induce isomorphisms between the bundles $J$ and $G_{i}$. This construction goes back to Labourie [40], who studied immersions of surfaces into hyperbolic 3-manifolds. The maps $K_{i}$ determine the isomorphisms $\mathcal{K}_{i}:\mathcal{A}\rightarrow\mathcal{A}$. (By isomorphisms between Banach spaces we mean bounded linear bijections with bounded inverse.) The metric $s$ similarly defines the maps $K$ and $\mathcal{K}$. The operators $\mathcal{K}_{i}$ converge to $\mathcal{K}$ in the sense of operator norms.
The tangent space to $C^{\infty}(J)$ at $j^{\prime}$ is defined as the space of smooth operator fields $z^{\prime}:TS\rightarrow TS$ satisfying $z^{\prime}j^{\prime}+j^{\prime}z^{\prime}=0$. Denote the respective bundle by $Z^{\prime}$. Let $\nabla_{i}$ be the connections of $s_{i}$. The pullback of the linearization of $d^{\nabla_{i}}$ to $C^{\infty}(Z^{\prime})$ by $K_{i}$ is given by

### v2

@@ 3.1. Equivariant surfaces in Minkowski 3-space
$\langle a_{1},a_{2}\rangle_{L^{2}}=\int_{S}{\rm tr}(a_{1}a_{2})d{\rm area}_{\sigma}.$
The element $\zeta_{\sigma}(a)$ is orthogonal to the tangent space to the isotopy orbit of $\sigma$ if and only if $a$ is Codazzi for $\sigma$. See, e.g., [69, Section 3.2]. It is also easy to notice that $\zeta_{\sigma}(a)$ is orthogonal to the tangent space to the conformal orbit of $\sigma$ if and only if $a$ is traceless. We arrive to the well-known fact that that the restriction of $\eta_{\sigma}$ to the subspace of $\sigma$-symmetric traceless Codazzi operator fields is an isomorphism. When $\sigma$ is the hyperbolic representative of the conformal class, the $L^{2}$-scalar product produces the Weil–Petersson scalar product on the Teichmüller space, up to a universal constant depending on a convention. It also follows from the discussion that if $x_{1},x_{2}\in T_{\sigma}\mathscr{T}$ and $a_{1}\in\eta_{\sigma}^{-1}(x_{1})$, $a_{2}\in\eta_{\sigma}^{-1}(x_{2})$ are such that $a_{1}$ and $a_{2}$ are Codazzi and $a_{2}$ is traceless, then
$\langle x_{1},x_{2}\rangle_{WP}=\int_{S}{\rm tr}(a_{1}a_{2})d{\rm area}_{\sigma}.$
Now return to an equivariant isometric immersion $(\rho,f)$ with induced metric $s$ and shape operator $b$. Let $\rho_{F}$ be the linear part of $\rho$ and $\sigma$ be the hyperbolic metric on $S$ given by $\mathbb{H}^{2}/\rho_{F}$ and pulled to $S$ by the composition of the Gauss map from f(S~)f($\widetilde${S}) and of $f$. Define $a:=b^{-1}$, i.e., $a={\rm inv}\circ(\chi_{s,+}^{0})^{-1}(\sigma)$. Then $s=\sigma(a\cdot,a\cdot)$ and $a$ is Codazzi for $\sigma$, see, e.g., [16, Proposition 2.7] of Bonsante–Seppi. Define
$F(s,\sigma):=\int_{S}{\rm tr}(a)d{\rm area}_{\sigma}.$
Due to Theorem 3.2, $F$ is indeed a well-defined function of a negatively curved metric $s$ and a hyperbolic metric $\sigma$. Furthermore, it depends only on the isotopy classes of $s$, $\sigma$. For fixed $s$, consider $F(s,\cdot)$ as a function $F_{s}$ over the Teichmüller space $\mathscr{T}$. It was shown by Smith in [62, Lemma 5.3] that
@@
Pick $x\in T_{\sigma}\mathscr{T}$. If $a^{\prime}\in\eta_{\sigma}^{-1}(x)$ is traceless and Codazzi, then
$\dot{F}_{s}=-\int_{S}{\rm tr}(aa^{\prime})d{\rm area}_{\sigma}.$
@@
The vector field $x^{\prime}_{s}$ is the Weil–Petersson gradient field of $F_{s}$ and $x_{s}$ is the symplectic Weil–Petersson gradient field of $F_{s}$.
Furthermore, Smith proved [62, Lemma 5.6]
@@
For any two negatively curved metrics $s_{+}$ and $s_{-}$, the vector fields $x_{s_{+}}^{\prime}$ and $x_{s_{-}}^{\prime}$ intersect transversely at every intersection point as submanifolds of $T\mathscr{T}$. The same holds for $x_{s_{+}}$ and $x_{s_{-}}$.
@@
If at a point $\sigma\in\mathscr{T}$ we have $x_{s_{+}}^{\prime}(\sigma)=-x_{s_{-}}^{\prime}(\sigma)$, then the Weil–Petersson gradients of functions $F_{s_{+}}$ and $F_{s_{-}}$ coincide at $\sigma$. Thus, $\sigma$ is a critical point of $F_{s_{+}}+F_{s_{-}}$. Since the latter function is strictly convex with respect to the Weil–Petersson metric, its Weil–Petersson Hessian is non-degenerate. Thereby, the intersection point of $x_{s_{+}}^{\prime}$ and $x_{s_{-}}^{\prime}$ is transverse at $x_{s_{+}}^{\prime}(\sigma)=-x_{s_{-}}^{\prime}(\sigma)$. The same holds for the vector fields $x_{s_{+}}$, $x_{s_{-}}$. ∎
@@
[…]
@@
[…]
@@
Due to Theorem 3.2, expression (3.3) is nonzero. But this contradicts Corollary 3.6, as the images of $\phi_{s_{+},+}$ and $\phi_{s_{-},-}$ are exactly the vector fields $x_{s_{+}}$ and $x_{s_{-}}$, when the latter are considered as submanifolds of $T\mathscr{T}$. ∎
@@ 3.2. The infinitesimal Pogorelov map
Recall the definition of $\mathbb{A}d\mathbb{S}^{3}$ from Section 2.1. We now pick a convenient affine chart for it, given by the projection to the plane $\{x\in\mathbb{R}^{2,2}:x_{4}=1\}$. This plane is also a model for Minkowski 3-space, and we will consider it endowed simultaneously with both geometries. We denote it now by $\mathbb{R}^{2,1}$, and by the origin we will mean the point $o=(0,0,0,1)$. The part of $\mathbb{A}d\mathbb{S}^{3}$ that is captured in this plane is the interior of a one-sheeted hyperboloid
@@
Here, for convenience, we only consider the restrictions of the Killing fields to $C$, though one can define $\Phi$ over the whole tangent bundle $TA$ of $A$, see, e.g., [22]. See [26, Lemma 11] of Fillastre for a proof of Lemma 3.8, which is based on [57, Lemma 1.10] of Schlenker. In particular, $\Phi$ determines an isomorphism $\Psi:\mathfrak{g}\rightarrow\mathfrak{g}^{0}$ of Lie algebras, considered as vector spaces. Furthermore, every $\gamma\in G_{F}$ preserves the decomposition into radial and lateral parts, as well as the function $r$. Hence, $\Phi$ is $G_{F}$-equivariant. In turn, this implies
@@
For a Fuchsian $\rho:\pi_{1}S\rightarrow G_{F}$, Lemma 3.9 establishes a canonical isomorphism
$\Psi_{\rho}:Z^{1}_{\rho}(\mathfrak{g})\rightarrow Z^{1}_{\rho}(\mathfrak{g}^{0}),$
which sends coboundaries onto coboundaries, and hence also induces an isomorphism of cohomologies. Now we also note
@@
Let $\rho\in\widehat{\mathcal{R}}_{F}$, $\Sigma\subset C$ be a $\rho$-invariant submanifold and $v$ be a vector field over $\Sigma$, automorphic with respect to $\dot{\rho}\in Z^{1}_{\rho}(\mathfrak{g})$. Then $\Phi(v)$ is automorphic with respect to $\Psi_{\rho}(\dot{\rho})\in Z^{1}_{\rho}(\mathfrak{g}^{0})$.
@@
$d\rho(\gamma)\big(\Phi(v(x))\big)+\Psi_{\rho}(\dot{\rho})(\gamma)(\rho(\gamma)x)=\frac{d\rho(\gamma)(v_{r}(x))+\dot{\rho}(\gamma)_{r}(\rho(\gamma)x)}{1-r^{2}(x)}+$
$d\rho(\gamma)(v_{l}(x))+\dot{\rho}(\gamma)_{l}(\rho(\gamma)x)=\frac{v_{r}(\rho(\gamma)x)}{1-r^{2}(\rho(\gamma)x)}+v_{l}(\rho(\gamma)x)=\Phi(v(\rho(\gamma)x)),$
@@
Proof of Proposition 3.1..
[…]
@@
It is interesting to notice that even though one can define the infinitesimal Pogorelov map outside of $C$, this proof works only when the pair of embeddings is Fuchsian. The reason is that for a general equivariant embedding, its image in the projective model will not be equivariant with respect to some Minkowski representation. There is a nice way to do a rescaling, to obtain quasi-Fuchsian Minkowski spacetimes from anti-de Sitter ones (more precisely, from the components of the complements to the convex core), well-behaved on holonomies, see [8] by Benedetti–Bonsante. However, this is not a projective map, and hence does not have a related infinitesimal Pogorelov map. (An infinitesimal Pogorelov map can be associated to any projective map between ambient spaces, see [27]).
@@ 4.1. A bit of functional analysis
For a bounded operator $\mathcal{L}$ between two Banach spaces $\mathcal{A},\mathcal{B}$, we define its reduced minimum modulus
$\gamma(\mathcal{L}):=\inf\{\|\mathcal{L}a\|:a\in\mathcal{A},d(a,{\rm Ker}\mathcal{L})=1\}.$
When the range of $\mathcal{L}$ is closed, $\gamma(\mathcal{L})>0$. See, e.g., [38, Chapter IV, Theorem 5.2].
For two subspaces $A,B$ of a Banach space, we define their distance as
@@
Let $A$ be a finite-dimensional subspace in a Banach space, and $A_{i}$ be a sequence of subspaces, converging to $A$ in the distance (4.1). Let $a_{i}\in A_{i}$ be a uniformly bounded sequence of vectors. Then, up to subsequence, $a_{i}$ converges to $a\in A$.
@@
Since $A$ is finite-dimensional, for every $a_{i}$ there is $a^{\prime}_{i}\in A$ realizing the distance from $a_{i}$ to $A$. The sequence $a^{\prime}_{i}$ is uniformly bounded. Since bounded subsets of $A$ are precompact, a subsequence of $a^{\prime}_{i}$ converges to some $a$. It is easy to see that the same subsequence of $a_{i}$ also converges to $a$. ∎
@@ 4.2. Local infinitesimal rigidity
Proposition 4.2.
Let $e\in\mathcal{E}_{F}$. There exists a neighborhood $U_{e}$ of $e$ in $\mathcal{E}$ such that for every $e^{\prime}\in U_{e}$ the differential $d_{e^{\prime}}\mathcal{I}$ is injective.
@@
Consider $e^{\prime}\in\mathcal{E}$. Let $_{+}, ss_{-}$, $b_{+}$ and $b_{-}$ be the respective induced metrics and shape operators on $S$ given by $e^{\prime}$. The differential $d_{e^{\prime}}\mathcal{I}$ is injective if and only if the images of $\phi_{s_{+},+}$ and $\phi_{s_{-},-}$ intersect transversely at $\phi_{s_{+},+}(b_{+})=\phi_{s_{-},-}(b_{-})$.
@@
[…]
Now we need to show that if $s_{i}$ converge to $s$ and $\rho_{i}\in{\rm Im}(\phi_{s_{i},\pm})$ converge to $\rho\in{\rm Im}(\phi_{s,\pm})$, then the tangent space to ${\rm Im}(\phi_{s_{i},\pm})$ at $\rho_{i}$ converge to the tangent space to ${\rm Im}(\phi_{s,\pm})$ at $\rho$. Let $\mathcal{A}$ be the Sobolev space $W^{1,2}(T^{1}_{1}S)$ of sections of $T^{1}_{1}S$ that have finite $L^{2}$-norm and so do their first weak derivatives. We consider $\mathcal{A}$ as a Banach space. The desired convergence result follows from
@@
For every $i$ denote by $G_{i}\subset T^{1}_{1}S$ the subbundle of $s_{i}$-symmetric operator fields $TS\rightarrow TS$ satisfying the Gauss equation for $s_{i}$. Denote by $J\subset T^{1}_{1}S$ the subbundle of almost-complex structures on $S$ (compatible with the orientation) and let $_{i}, jj$ be the almost-complex structures of $_{i}, ss$. Consider the bundle isomorphisms $K_{i}:T^{1}_{1}S\rightarrow T^{1}_{1}S$ given by
@@
where $\lambda_{i}:=\sqrt{-\kappa_{i}-1}$ for the curvature $\kappa_{i}$ of $s_{i}$. Notice that $K_{i}$ induce isomorphisms between the bundles $J$ and $G_{i}$. This construction goes back to Labourie [42], who studied immersions of surfaces into hyperbolic 3-manifolds. The maps $K_{i}$ determine the isomorphisms $\mathcal{K}_{i}:\mathcal{A}\rightarrow\mathcal{A}$. (By isomorphisms between Banach spaces we mean bounded linear bijections with bounded inverse.) The metric $s$ similarly defines the maps $K$ and $\mathcal{K}$. The operators $\mathcal{K}_{i}$ converge to $\mathcal{K}$ in the sense of operator norms. Define $a_{i}:=\mathcal{K}_{i}^{-1}b_{i}$, $a:=\mathcal{K}^{-1}b$. Then $a_{i}$ $C^{\infty}$-converge to $a$.
The tangent space to $C^{\infty}(J)$ at arbitrary $j^{\prime}\in C^{\infty}(J)$ is defined as the space of smooth operator fields $z^{\prime}:TS\rightarrow TS$ satisfying $z^{\prime}j^{\prime}+j^{\prime}z^{\prime}=0$. Denote the respective bundle by $Z^{\prime}$. Let $\nabla_{i}$ be the Levi-Civita connections of $s_{i}$. The pullback of the linearization of $d^{\nabla_{i}}$ to $C^{\infty}(Z^{\prime})$ by $K_{i}$ is given by